# Kardia-R1: Unleashing LLMs to Reason toward Understanding and Empathy for Emotional Support via Rubric-as-Judge Reinforcement Learning

### v1

@@ 2.1 Empathetic Dialogue Generation
Large language models shift empathetic dialogue toward a reaso
ning-centered paradigm wei2022chain; guo2025deepseek, in which instruction-tuned assistants use counseling-style multi-turn data and structured prompts to analyze user experiences before generating supportive responses chen2023soulchat. Prompt-engineering approaches further embed appraisal
-theoretic or commonsense-driven reasoning hu2024aptness; dai2025psyche; cai2024empcrl, encouraging models to infer emotional drivers, evaluate situational appraisals, articulate multi-step empathetic cognition, and calibrate emotional strategies beyond generic comfort. Additional methods synthesize affective signals—such as emotional causes, user intent, and personality traits he2025ecc; xie2021empathetic; yuan2025reflectdiffu; wu2025traits—and incorporate them into finetuning bi2023diffusemp or reinforcement-learning pipelines li2024reinforcement; sotolar2024empo; dai2025psyche via embedding-based rewards on EmpatheticDialogues rashkin2019towards. Although such LLM-based emotional support agents demonstrate strong cognitive-empathy reasoning and stable affective language xu2025multiagentesc, they still struggle to sustain user-specific alignment and safe long-horizon behavior as needs evolve, and they incur high inference and adaptation costs when required to cover increasingly diverse emotional scenarios. We address these inefficiencies by training LLMs to internalize a perspective-taking empathy that integrates user comprehension, contextual reasoning, emotion recognition, and empathetic response generation, thereby improving generalization to new users and shifting emotional contexts.
@@ 4 Kardia-R1
We first perform cold-start supervised fine-tuning of a base LLM $\pi_{\theta_{0}}$ on emotionally simple trajectories $\mathcal{D}{\mathrm{easy}}$ using context-response pairs, where $c_{t}$ includes user identity, emotional state, and dialogue history up to turn $t$, and $y_{t}$ is formatted into four explicit spans delineated by <|understanding_begin|><|understanding_end|>, <|reasoning_begin|> <|reasoning_end|>, <|emotion_begin|><|emotion_end|>, and <|response_begin|> <|response_end|> tags. This structured representation scaffolds model learning by disentangling key empathetic components while preserving their generation order. The model is trained to maximize the log-likelihood of the full target sequence under a standard autoregressive objective:
@@
where we adpot equal weighting ($\lambda_{f}=\lambda_{e}=\lambda_{r}=\tfrac{1}{3}$) to ensure a balanced contribution from structural accuracy, affective alignment, and rubric-based communicative quality. 1. Format reward. We utilize a format reward to enforce strict adherence to the required four-span structure for progressive understanding, reasoning, emotion identification, and empathetic response generation. For a candidate $o_{j}$, the model must produce four spans delimited by <|understanding_begin|><|understanding_end|>, <|reasoning_begin|> <|reasoning_end|>, <|emotion_begin|> <|emotion_end|>, and <|response_begin|> <|response_end|>. Format checking evaluates (i) the presence of all begin/end tags, (ii) the ordering of spans, and (iii) non-empty content within each segment. Let $v_{j,k}\in\{0,1\}$ indicate violation of the $k$-th constraint for candidate $j$, with $K$ denoting the total number of constraints, formally:

### v2

@@ 2.1 Empathetic Dialogue Generation
Large language models shift empathetic dialogue toward a reasoning-centered paradigm wei2022chain; guo2025deepseek, in which instruction-tuned assistants use counseling-style multi-turn data and structured prompts to analyze user experiences before generating supportive responses chen2023soulchat. Prompt-engineering approaches further embed appraisal-theoretic or commonsense-driven reasoning hu2024aptness; dai2025psyche; cai2024empcrl, encouraging models to infer emotional drivers, evaluate situational appraisals, articulate multi-step empathetic cognition, and calibrate emotional strategies beyond generic comfort. Additional methods synthesize affective signals—such as emotional causes, user intent, and personality traits he2025ecc; xie2021empathetic; yuan2025reflectdiffu; wu2025traits—and incorporate them into finetuning bi2023diffusemp or reinforcement-learning pipelines li2024reinforcement; sotolar2024empo; dai2025psyche via embedding-based rewards on EmpatheticDialogues rashkin2019towards. Although such LLM-based emotional support agents demonstrate strong cognitive-empathy reasoning and stable affective language xu2025multiagentesc, they still struggle to sustain user-specific alignment and safe long-horizon behavior as needs evolve, and they incur high inference and adaptation costs when required to cover increasingly diverse emotional scenarios. We address these inefficiencies by training LLMs to internalize a perspective-taking empathy that integrates user comprehension, contextual reasoning, emotion recognition, and empathetic response generation, thereby improving generalization to new users and shifting emotional contexts.
@@ 4 Kardia-R1
We first perform cold-start supervised fine-tuning of a base LLM $\pi_{\theta_{0}}$ on emotionally simple trajectories $\mathcal{D}{\mathrm{easy}}$ using context-response pairs, where $c_{t}$ includes user identity, emotional state, and dialogue history up to turn $t$, and $y_{t}$ is formatted into four explicit spans delineated by <|understanding_begin|><|understanding_end|><|understanding\_begin|><|understanding\_end|>, <|reasoning_begin|><|reasoning_end|><|reasoning\_begin|><|reasoning\_end|>, <|emotion_begin|><|emotion_end|><|emotion\_begin|><|emotion\_end|>, and <|response_begin|><|response_end|><|response\_begin|><|response\_end|> tags. This structured representation scaffolds model learning by disentangling key empathetic components while preserving their generation order. The model is trained to maximize the log-likelihood of the full target sequence under a standard autoregressive objective:
@@
where we adpot equal weighting ($\lambda_{f}=\lambda_{e}=\lambda_{r}=\tfrac{1}{3}$) to ensure a balanced contribution from structural accuracy, affective alignment, and rubric-based communicative quality. 1. Format reward. We utilize a format reward to enforce strict adherence to the required four-span structure for progressive understanding, reasoning, emotion identification, and empathetic response generation. For a candidate $o_{j}$, the model must produce four spans delimited by <|understanding_begin|><|understanding_end|><|understanding\_begin|><|understanding\_end|>, <|reasoning_begin|><|reasoning_end|><|reasoning\_begin|><|reasoning\_end|>, <|emotion_begin|><|emotion_end|><|emotion\_begin|><|emotion\_end|>, and <|response_begin|><|response_end|><|response\_begin|><|response\_end|>. Format checking evaluates (i) the presence of all begin/end tags, (ii) the ordering of spans, and (iii) non-empty content within each segment. Let $v_{j,k}\in\{0,1\}$ indicate violation of the $k$-th constraint for candidate $j$, with $K$ denoting the total number of constraints, formally: